

********************************************************************
UARS WEEKLY STATUS REPORT - 14 APR 2000
********************************************************************
***********
SOLSTICE
***********
We continue to gather enough data to meet all the SOLSTICE science
objectives during this period of reduced tape recorder operations.
SOLSTICE operates normally, collecting solar data on all available
orbital tracking periods and stellar calibration observations whenever
possible. Level 2 processing is halted due to bugs in the OA_SSPP_ATT
service. CDHF and CSC personnel are working on the problem.
***********
SUSIM
***********
SUSIM continues to gather a near normal set of daily solar scans
scheduled for periods of predicted telemetry.
Last week, this report mentioned that SUSIM did not take data on
UARS day 93. This date was incorrect. It should have read
"UARS day 97", instead.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
The HRDI instrument is operating normally, collecting data on every
daytime portion of an orbit.
The HRDI operations team is completing an upgrade of the flight modes
which compensates for some of the mechanical aging by operating the
instrument in a more closed loop fashion, i.e. instrument timing is based
on actual positions of various components (telescope position, filter wheel
position) rather than conservative timing estimates.
Work proceeds on a new data delivery, which will allow the data to be
processed properly with the frequent dropouts. In addition, algorithm
refinements and a new product, mesospheric ozone, are being added.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues with no
problems. The MPG continues to assist the Project Scientistwith the scheduling
of instrument operations.
***********
CDHF
***********
STATUS
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
- All requested reprocessing is complete as of this reporting period.
UARS CDHF REHOST:
- Successful processing on Rehost System:
Attitude Extraction, Observatory Log, Sensor Extraction, HALOE, MLS,
SUSIM, Solstice, and WINDII.
- MLS V4.22 (OUTPUT 4) has tested successfully on Rehost System. MLS
has validated the output product. Thanks for the timely response.
- Preparing to test ISAMS V12.0 and PEM V3.32 software on the
Rehost System.
***********
FOT
***********
REPORTING PERIOD
This report covers 01 April, 2000 (Orbit 46759, GMT Day 092) through 07 April
2000 (Orbit 46862, GMT Day 098).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 096/2255 46832 AXIS Off
WINDII 097/0036 46833 On
The beta angle ranged between the angles of 38.1 degrees to 49.2 degrees for
this report period. The beta angle is now increasing toward a maximum angle of
+49.0 degrees on 08 April 2000 (DOY 099).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 0.9 to 1.1 degrees C. Additional battery performance data for each
battery on 17 March 2000 and 07 April 2000 are provided in Appendix D.
The clock error ranged from 11.2 to -11.1 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
% OUT-OF-LIMITS = 01
% ALERTS = 02
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
% FOT = 20
% HRDI = 04
% PEM = 01
% WINDII = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
% FOT = 05
Five (5) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
- Transponder A: 00/091 - 2287.496948
- Transponder B: 00/091 - 2287.499113
GROUND SYSTEM OPERATION
Five (5) operational support problem occurred this period. Anomaly Reports are
included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 00/097 show the observatory data loss to be 301 hours, 12
minutes, 33.513 seconds* (an increase of 0 minutes 0.00 seconds since the last
report period).
*These calculations have been recalculated to include 1999/346 through 2000/006
statistics using scheduled times.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT corrected the halt of the ATCP (refer to AIR-001/00) on 03 April 00.
The halt was caused by the CMS incorrectly reporting the uplink time for the
new load. Thus, the load was sent before it should have been, which overwrote
the active ATC load. The location overwritten was overwritten by and ATC halt
command and that command went out. This was not a case of operator error,
however, because the CMS had reported a load after time of 17:53:11 GMT when it
really wasn't safe to uplink the new load until after 18:01:13 GMT.
In order to avoid this problem in the future the ATC uplink times will be
checked manually to make sure they are valid. A DR will also be written so
that a fix to the CMS can be made that will perform this check automatically.
There were more problems encountered with the UTTS this week, but all have been
fixed and the FOT is ready to begin testing of the new TMONs next week.
The FOT also began developing an online procedure for safehold should it happen
before the new TMONs are in place or during and OBC halt. This procedure has
yet to be finalized and must undergo peer review before it will be ready.
MEETINGS
The FOT attended the UARS monthly status meeting 05 April 00.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for April 25, 2000 (DOY 116). This
maneuver will be a Reverse to Forward Flight maneuver (Northern Hemisphere
viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 07 April 2000
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
46760 092/0350 00-132 NBTR NBTR OPS DAY-092
46761 092/0506 00-132 NBTR NBTR OPS DAY-092
46769 092/1759 00-135 NBTR NBTR OPS DAY-093
46771 092/2207 00-136 OBC Clk Adj 0.00 to 20.6 msec/day
46776 093/0542 00-135 NBTR NBTR OPS DAY-093
46784 093/1757 00-137 NBTR NBTR OPS DAY-094
46799 094/1751 00-138 OBC Clk Adj 0.00 to 20.6 msec/day
46799 094/1751 00-139 NBTR NBTR OPS
46799 094/1751 00-140 HRDI HRDI MICROLOAD
46799 094/1851 00-141 OBC ATC POINTER
46799 094/1851 00-142 HGA HGA SLEW
46803 094/2356 00-143 HRDI CHECK TABLE INTERPRETER
46803 095/0203 00-143 HRDI CHECK TABLE INTERPRETER
46805 095/0401 00-139 NBTR NBTR OPS
46806 095/0513 00-139 NBTR NBTR OPS
46813 095/1713 00-144 HRDI HRDI MICROLOAD
46814 095/1854 00-145 NBTR NBTR OPS
46821 096/0553 00-145 NBTR NBTR OPS DAY-096
46830 096/2036 00-146 NBTR NBTR OPS DAY-097
46832 096/2255 00-147 AXIS-1 Power OFF
46833 097/0036 00-148 WINDII Power ON
46833 097/0036 00-149 OBC Clk Adj 0.00 to 20.6 msec/day
46835 097/0410 00-145 NBTR NBTR OPS DAY-097
46841 097/1240 00-149 NBTR NBTR OPS DAY-098
46852 098/0658 00-149 NBTR NBTR OPS DAY-098
46859 098/1802 00-150 NBTR NBTR OPS DAY-099
_______________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 07 April 2000
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HRPSEQBIMON Y-HI PSE QUIET BUS CURRENT MON 46806
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
SSCMDERR SSPP CMD ERROR 46797
HGSWLIMFAIL HGA SW LIMIT ERROR 46799
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 07 April 2000
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
46771 092/22:13 (01 Apr 00) from 0.00 to 20.6 msec/day
46799 094/17:53 (03 Apr 00) from 20.6 to 0.00 msec/day
46833 097/00:56 (06 Apr 00) from 0.00 to 20.6 msec/day
_______________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 07 April 2000
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
Battery performance data not available for this report period.
07 April 2000 - GMT Day 098 Beta = 49.2 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- -------- ---------- ------- ------- ------
1 -33.6/-56.0 -3.05/-4.31 +0.0/+0.0 15.8 0.000 0.0
2 +56.0/-5.6 +5.47/+4.19 +26.4/-16.0 24.6 1.042 21.3
3 +22.4/-22.4 +4.51/+3.23 +26.0/-16.0 24.5 1.040 21.0
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ------- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 07 April 2000
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-008 FOT 46799 094/1751 ATC Halt
00-005 FOT 46799 094/1851 14170 Generic Late Acq.
00-009 FOT 46813 095/1713 No File Xfer Cap.
00-010 FOT 46827 096/1428 No File Xfer Cap.
00-007 FOT 46845 097/2038 14170 Generic Late Acq.
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-005 FOT see list March 00 14170 Generic Late Acq.
00-008 FOT 46799 094/1751 ATC Halt
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-008 FOT 46799 094/1751 ATC Halt
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-064 FOT 44135 10/8-9/99 N/A HRDI Cal Lamp Errors
99-071 FOT 44398 299/1212 21512 TPF Not Receiv 512
00-007 FOT see list April 00 14170 Generic Late Acq.
00-009 FOT 46813 095/1713 No File Xfer Cap.
00-010 FOT 46827 096/1428 No File Xfer Cap.
SEVNTFW00/14